OpenStudy (anonymous):

What is the equation of a line that passes through the points (-3,6) and (1,2) ?

OpenStudy (anonymous):

slope-intercept equation of a line: y = mx + b, where m is slope and b is y-intercept. Slope is change in y over change in x: (2 - 6) / (1 - -3) = -1. Then you plug that in for m in the above equation and use one of the points given and plug that in for x and y (I'll use the second point): y = mx + b 2 = -1 + b b = 3 So the equation is y = -x + 3 Here's the graph: http://bit.ly/qOTgsn

OpenStudy (anonymous):

Why does b= 3 ?

OpenStudy (anonymous):

I plugged "2" from the second point in for y, "1" for x, and "-1" for m. Then I solved for b: y = mx + b (2) = (-1)(1) + b 2 = -1 + b Then add 1 to each side: 3 = b

OpenStudy (anonymous):

Can I use first point of y or is it necessary that I use second y to solve this equation?

OpenStudy (anonymous):

you can choose either point, because by definition they both satisfy the equation. I chose the second one because it looked easier.

OpenStudy (anonymous):

Oh i get it now tysm :)

OpenStudy (anonymous):

ymtw